A person travels by car from one city to another with different constant speeds between pairs of cities. She drives for 45.0 min at 50.0 km/h, 7.0 min at 80.0 km/h, and 50.0 min at 55.0 km/h and spends 40.0 min eating lunch and buying gas.

Average speed?

Distance between initial and final cities?
Answer #1

d1 = (45/60)* 50 = 37.5 km

d2 = (7/60)* 80 = 9.33 km

d3 = (50/60)* 55 = 45.833 km

======

a)

av speed = (d1 + d2 + d3) / ( t1 + t2 + t3 + t4)

s = (37.5 + 9.33 + 45.833) / ( 45/60 + 7/60 + 50/60 + 40/60)

s = 39.15 km/h

=========

b)

d = d1 + d2 + d3 = 92.663 m

=======

do comment in case any doubt, will reply for sure.. Goodluck